# MUGSQA: Novel Multi-Uncertainty-Based Gaussian Splatting Quality Assessment Method, Dataset, and Benchmarks

### v1

@@ Abstract
Gaussian Splatting (GS) has recently emerged as a promising technique for 3D object reconstruction, delivering high-quality rendering results with significantly improved reconstruction speed. As variants continue to appear, assessing the perceptual quality of 3D objects reconstructed with different GS-based methods remains an open challenge. To address this issue, we first propose a unified multi-distance subjective quality assessment method that closely mimics human viewing behavior for objects reconstructed with GS-based methods in actual applications, thereby better collecting perceptual experiences. Based on it, we also construct a novel GS quality assessment dataset named MUGSQA, which is constructed considering multiple uncertainties of the input data. These uncertainties include the quantity and resolution of input views, the view distance, and the accuracy of the initial point cloud. Moreover, we construct two benchmarks: one to evaluate the robustness of various GS-based reconstruction methods under multiple uncertainties, and the other to evaluate the performance of existing quality assessment metrics. Our dataset and benchmark code will be released soon.
@@ 1 Introduction
3D reconstruction is a fundamental problem in computer vision and computer graphics, aiming to recover accurate geometry and appearance of real-world objects and scenes. Among emerging approaches, the first method based on Gaussian Splatting (GS) [1] offers a compelling balance between high rendering quality and real-time performance, especially for 3D object reconstruction. Its outstanding performance quickly makes it one of the most promising solutions for practical deployment in 3D object reconstruction and draws attention from both academia and industry.
@@
Besides, existing Subjective Quality Assessment (SQA) methods often present the 3D object to the subjects with a fixed view or a single-distance display [18, 19], making it difficult to reflect the behavior of subjects when dynamically observing Gaussian objects [22, 23] in interactive or immersive scenarios. In order to better align with the above, we propose a unified multi-distance SQA method that guides observers to examine Gaussian objects from various distances and multiple views. Based on this, we conduct a large-scale subjective experiment to collect quality scores for the MUGSQA dataset. We gather 2,452 participants and ultimately obtain over 226,800 valid scores, ensuring that the scores we finally collect are sufficient and reliable. Finally, we construct two benchmarks based on the MUGSQA dataset to evaluate the robustness of GS-based reconstruction methods and the performance of existing objective quality assessment metrics on Gaussian objects. This fills the gap in the current evaluation system in this field and promotes the standardized development of GSQA.
@@ 2 Data Preparation
Additional Set. Next, we construct an additional set using more reconstruction methods. For this set, we only use 3 out of the 55 source models, but employ 5 GS-based methods for reconstruction: 3DGS [1], Mip-Splatting [3], Scaffold-GS [4], EAGLES [5], and Octree-GS [6]. All other settings remain consistent with the main set. Similarly, we filter this set and obtain 444 additional set samples. In total, we obtain $1,970+444=2,414$ reconstructed models. Figure 2 shows the overall data generation pipeline.
@@ 3 Subjective Quality Assessment
Experiment. To obtain reliable and controllable results, we start a crowdsourced project using MTurk  [^3] https://www.mturk.com and create a scoring interface. As shown in Figure 3, the interface includes three modules: reference video, distortion video, and scoring area. After each pair of videos is played, workers are allowed to slide the scoring bar. In the training stage, a suggested score and a reason corresponding to the distortion will be displayed. After training, participants can enter the test stage of the experiment, during which the suggested scores will no longer be displayed, and the rest of the content remains the same as in the training stage. At the end of the experiment, the scoring results will be automatically uploaded, and after our review, the participants will be paid. Ultimately, 226,800 quality scores are collected and a total of 2,452 participants complete the experiment.
@@ 4.1 Dataset Completion and Comparison
As a result, we retain 101,555 valid scores, ensuring that each sample in every playlist has at least 30 valid scores. Then we compute Mean Opinion Scores (MOS) by averaging the scores given by different participants on each stimulus. Similarly to [17], we map the MOS to a continuous range of 0 to 10, where higher scores represent better quality. At this point, we have completed the dataset.
As shown in Table 1, our dataset has several advantages over existing datasets. Firstly, MUGSQA compensates for the deficiencies in GT by using synthetic data. This not only includes the image data required for reconstruction, but also contains the 3D mesh models, providing more reliable comparisons and analyses. Secondly, MUGSQA addresses the shortcomings of existing datasets in single-object reconstruction. Most datasets only contain scenes, whereas our dataset encompasses 55 synthetic objects as source models. In fact, if a single object can be reconstructed, it will be more conducive to an in-depth analysis of the distortion characteristics and metric design. This need to assess the quality of a single Gaussian object is crucial in scenarios requiring a large number of high-quality synthetic objects [29, 30]. In terms of SQA methods, compared to other datasets that only render frames in a fixed scale, our dataset takes into account the quality differences generated by rendering at different scales, thereby achieving 180 rendered frames and covering as many as 3 cycles. In terms of data annotation, our subjective experiments are also more thorough, including 2,414 valid MOS.
@@ 4.2 Robustness of GS-based Reconstruction Methods
To evaluate the robustness of different GS-based reconstruction methods using the MUGSQA dataset, we define a robustness score $R_{u}\in[0,100]$, which integrates three aspects: stability, consistency, and performance. Stability is derived from the coefficient of variation $CV=\frac{\sigma}{\mu}\times 100\%$, consistency from the MOS range $M=\max_{i}\{MOS_{i}\}-\min_{i}\{MOS_{i}\}$, and performance from the mean MOS $\mu$. These are mapped to $[0,100]$ and combined as:
$\begin{split}R_{u}=\ &0.4\times\max(0,\ 100-2\times CV)\ +\\ &0.3\times\max(0,\ 100-20\times M)\ +\\ &0.3\times\min(100,\ 10\times\mu)\end{split}.$ (2)

### v3

@@ Abstract
Gaussian Splatting (GS) has recently emerged as a promising technique for 3D object reconstruction, delivering high-quality rendering results with significantly improved reconstruction speed. As variants continue to appear, assessing the perceptual quality of 3D objects reconstructed with different GS-based methods remains an open challenge. To address this issue, we first propose a unified multi-distance subjective quality assessment method that closely mimics human viewing behavior for objects reconstructed with GS-based methods in actual applications, thereby better collecting perceptual experiences. Based on it, we also construct a novel GS quality assessment dataset named MUGSQA, which is constructed considering multiple uncertainties of the input data. These uncertainties include the quantity and resolution of input views, the view distance, and the accuracy of the initial point cloud. Moreover, we construct two benchmarks: one to evaluate the robustness of various GS-based reconstruction methods under multiple uncertainties, and the other to evaluate the performance of existing quality assessment metrics. Our dataset and code are available at https://github.com/Solivition/MUGSQA.
@@ 1 Introduction
3D reconstruction is a fundamental problem in computer vision, aiming to recover accurate geometry and appearance of real-world objects and scenes. Among emerging approaches, the first method based on Gaussian Splatting (GS) [7] offers a compelling balance between high rendering quality and real-time performance. Its outstanding performance quickly makes it one of the most promising solutions for practical deployment in 3D object reconstruction and draws attention from both academia and industry.
@@
Besides, existing Subjective Quality Assessment (SQA) methods often present the 3D object to the subjects with a fixed view or a single-distance display [30, 14], making it difficult to reflect the behavior of subjects when dynamically observing Gaussian objects [23] in interactive or immersive scenarios. In order to better align with the above, we propose a unified multi-distance SQA method that guides observers to examine Gaussian objects from various distances and multiple views. Based on this, we conduct a large-scale subjective experiment to collect quality scores for the MUGSQA dataset. We gather 2,452 participants and ultimately obtain over 226,800 valid scores, ensuring that the scores we finally collect are sufficient and reliable. Finally, we construct two benchmarks based on the MUGSQA dataset to evaluate the robustness of GS-based reconstruction methods and the performance of existing objective quality assessment metrics on Gaussian objects. This fills the gap in the current evaluation system in this field and promotes the standardized development of GSQA.
@@ 2 Data Preparation
Additional Set. Next, we construct an additional set using more reconstruction methods. For this set, we only use 3 out of the 55 source models, but employ 5 GS-based methods for reconstruction: 3DGS [7], Mip-Splatting [27], Scaffold-GS [11], EAGLES [3], and Octree-GS [17]. This choice is made to keep the subjective experiment manageable while still covering diverse geometric and textural characteristics. All other settings remain consistent with the main set. Similarly, we filter this set and obtain 444 additional set samples. In total, we obtain $1,970+444=2,414$ reconstructed models. Figure 2 shows the overall data generation pipeline.
@@ 3 Subjective Quality Assessment
Experiment. To obtain reliable and controllable results [6], we start a crowdsourced project using MTurk  [^3] https://www.mturk.com and create a scoring interface. As shown in Figure 3, the interface includes three modules: reference video, distortion video, and scoring area. After each pair of videos is played, workers are allowed to slide the scoring bar. In the training stage, a suggested score and a reason corresponding to the distortion will be displayed. After training, participants can enter the test stage of the experiment, during which the suggested scores will no longer be displayed, and the rest of the content remains the same as in the training stage. At the end of the experiment, the scoring results will be automatically uploaded, and after our review, the participants will be paid. Ultimately, 226,800 quality scores are collected and a total of 2,452 participants complete the experiment.
@@ 4.1 Dataset Completion and Comparison
As a result, we retain 101,555 valid scores, ensuring that each sample in every playlist has at least 30 valid scores. Then we compute Mean Opinion Scores (MOS) by averaging the scores given by different participants on each stimulus. Similarly to [25], we map the MOS to a continuous range of 0 to 5, where higher scores represent better quality. At this point, we have completed the dataset.
As shown in Table 1, our dataset has several advantages over existing datasets. Firstly, MUGSQA compensates for the deficiencies in GT by using synthetic data. This not only includes the image data required for reconstruction, but also contains the 3D mesh models, providing more reliable comparisons and analyses. Secondly, MUGSQA addresses the shortcomings of existing datasets in single-object reconstruction. Most datasets only contain scenes, whereas our dataset encompasses 55 synthetic objects as source models. In fact, if a single object can be reconstructed, it will be more conducive to an in-depth analysis of the distortion characteristics and metric design. This need to assess the quality of a single Gaussian object is crucial in scenarios requiring a large number of high-quality synthetic objects [10]. In terms of SQA methods, compared to other datasets that only render frames in a fixed scale, our dataset takes into account the quality differences generated by rendering at different scales, thereby achieving 180 rendered frames and covering as many as 3 cycles. In terms of data annotation, our subjective experiments are also more thorough, including 2,414 valid MOS.
@@ 4.2 Robustness of GS-based Reconstruction Methods
To evaluate the robustness of different GS-based reconstruction methods using the MUGSQA dataset, we define a robustness score $R_{u}\in[0,100]$, which integrates three aspects: stability, consistency, and performance [19]. Stability is derived from the coefficient of variation $CV=\frac{\sigma}{\mu}\times 100\%$, consistency from the MOS range $M=\max_{i}\{MOS_{i}\}-\min_{i}\{MOS_{i}\}$, and performance from the mean MOS $\mu$. These are mapped to $[0,100]$ and combined as:
$\begin{split}&R_{u}=\ 0.4\times\max(0,\ 100-2\times CV)\ +\\ &0.3\times\max(0,\ 100-20\times M)\ +0.3\times\min(100,\ 10\times\mu)\end{split}.$ (2)
@@ 5 Conclusion
6 Acknowledgement
This research is partially supported by the Ministry of Education, Singapore, under the funding of MOE-T2EP20123-0006.